OpenStudy (anonymous):

mr.johnson invested $4000, part at 5% and the remainder at 3%. the total annual income from both investment was $152. find the amount invested at each rate.

OpenStudy (anonymous):

I_1 + I_2 = $152 1_1 = 5% x I_2 = (4,000 - x) * 3% * 1 => 5% x + (4,000 - x) * 3% = $152

OpenStudy (anonymous):

Do you follow me?

OpenStudy (anonymous):

what is _

OpenStudy (anonymous):

Interest 1 and Interest 2

OpenStudy (anonymous):

375?

OpenStudy (anonymous):

2% x + $120 = $152

OpenStudy (anonymous):

Can you find x from here?

OpenStudy (anonymous):

1,500

OpenStudy (anonymous):

2%x = 152 - 120 = 32 =>x = 32 * 100/ 2 = .....

OpenStudy (anonymous):

Since x is amount invest in 5% Can you find the other 3% amount ?

OpenStudy (anonymous):

Have you finished it? If you have, then you should understand how to solve your next post!